# Visual Imitation Enables Contextual Humanoid Control

### v2

@@ 2 Related Work
Recent progress in legged-robot motor skills follows two complementary streams. Reward-based methods use model-free RL in simulation, shaping behavior with handcrafted objectives that mix task terms (e.g., velocity tracking) and motion-naturalness regularizers; thanks to massive parallel physics engines [11, 12], this paradigm has produced agile locomotion on quadrupeds and humanoids without motion data. However, each new behavior demands tuning of user-defined rewards and environment scripting [13, 14, 15, 4, 6, 9]. Data-driven methods instead imitate reference motion, originally MoCap clips or monocular video, training a simulated character to track them and porting the idea to robots [1, 3, 16, 17, 7, 8]. For example, recent work [18, 19] frames legged locomotion as a next-token prediction task and pre-trains a policy on human data in kinematic space, showing strong performance. While imitation bypasses reward engineering [5], existing works typically assume flat ground or manually designed setups, limiting context-aware whole-body control; even animation systems that model human-scene contact rely on instrumented MoCap stages and thus lack scalability [20, 21]. Our system conditions on visual observations, the local height‑maps, and learns environment‑aware skills such as stair‑climbing and chair‑sitting directly from monocular RGB videos. Joint 4D human–scene reconstruction provides physically consistent reference motions, which RL distills into policies that transfer to a real humanoid (Table 1).
@@ 3 Real-to-Sim Data Acquisition
Our real-to-sim pipeline proceeds as summarized in Figure 2. We extract per-frame human poses and a raw scene point cloud from the input video (Sec. 3.1); jointly optimize them to obtain metrically aligned human trajectories and scene geometry (Sec. 3.2); apply gravity alignment and convert the filtered point cloud into a lightweight mesh (Sec. 3.3); and retarget the refined trajectories to the humanoid under joint-limit, contact, and collision constraints. (Sec. A.3) The resulting motion–mesh pairs are ready for policy learning in Section 4.
@@ 4 Policy Learning
Stage 1: MoCap Pre-Training MoCap pre-training lets a policy learn challenging skills from noisy video reconstruction while keeping hand-crafted priors to a minimum and bridging the human-to-robot embodiment gap. Earlier work tackles this either by sampling start poses with multi-agent RL [3] or by having a privileged simulator imitate the motion [5]. Radosavovic et al.[19] instead employed a form of kinematic pre‑training on human data. We adopt a simpler yet effective strategy: first pre‑train the policy on MoCap trajectories, then fine‑tune it on our reconstructed video data—both stages use reinforcement learning in a physics simulator. Even the MoCap‑only policy can be deployed directly on the real robot as shown in Fig 9. We used LAFAN motion capture data [54] retargeted to Unitree G1. For the pretrained policies, the conditioning the policy receives is the target joint angles, target root roll/pitch, and desired root direction.
@@ 5.1 Reconstruction and Data
We compare our method against baselines following the standard evaluation protocol [57, 45]. As summarized in Tab. 2, our method consistently achieves the best performance, outperforming prior work in both human trajectory accuracy (WA/W-MPJPE) and scene geometry (Chamfer Distance). Refer to Sec. A.4 for more evaluation details.
@@ 5.2 Real-world Deployment
Setup. We deploy our controller on a 23-DoF Unitree G1 humanoid and run it onboard at 50 Hz. Following [58], we set relatively low joint gains, $K_{p}=75$, to avoid excessively fast or overly stiff behaviour—which helps to avoid excessively violent contact when the robot makes heavy contact with objects such as chairs or stairs. Height-maps are computed in real time using Fast-lio2 [59] and probabilistic terrain mapping [60], respectively. We feed joystick targets from a human operator. Including policy running, all operations are run onboard. We found two critical ingredients for successful motion deployment through iterative sim-to-real trials: (i) relaxing the episode-termination tolerances with respect to the reference motion, and (ii) injecting realistic physics perturbations during training. Full details are given in Sec. C.
@@ Acknowledgments
We thank Brent Yi for his guidance with the excellent 3D visualization tool we use, Viser. We are grateful to Ritvik Singh, Jason Liu, Ankur Handa, Ilija Radosavovic, Himanshu Gaurav-Singh, Haven Feng, and Kevin Zakka for helpful advice and discussions during the paper. We thank Lea Müller for helpful discussions at the start of the project. We thank Zhizheng Liu for helpful suggestions on evaluating human and scene reconstruction. We thank Eric Xu, Matthew Liu, Hayeon Jeong, Hyunjoo Lee, Jihoon Choi, Tyler Bonnen, and Yao Tang for their help in capturing and featuring in the video clips used in this project. We acknowledge support from the BAIR Humanoid Intelligence Center. Chung Min Kim is supported by the NSF Research Fellowship Program, Grant DGE 2146752. Pieter Abbeel holds concurrent appointments as a Professor at UC Berkeley and as an Amazon Scholar. This paper describes work performed at UC Berkeley and is not associated with Amazon. This project was funded in part by NSF: CNS-2235013, IARPA DOI/IBC No. 140D0423C0035, ONR MURI award N00014-21-1-2801, and Bakar fellows.
@@ A.2 Objectives for Joint Human–Scene Reconstruction
As shown in Fig. 7(a), the coarse initialization produces inaccurate human trajectories and an incorrect scene scale. Our optimization procedure refines these estimates by jointly optimizing the human’s global translations ($\gamma^{1:T}$), global orientations ($\phi^{1:T}$), local poses ($\theta^{1:T}$), and the scene point cloud scale ($\alpha$). The optimization aligns the human and world as shown in Fig. 7(b), and aids generating simulation-ready data.
@@ B.2 Observations
Table 3 details the actor and critic observations. We note that, unlike works in graphics [1, 3], the actor gets all unprivileged observations that are all available on the real robot (see Section 5). Furthermore, the target root reference during train time is fed into the global frame to ensure it is aligned with the terrain. Thus, unlike [5, 8], we are able to train policies using a global root reference, allowing our policies to correctly imitate references in the world frame over long horizons, which is critical for correctly imitating motions tied to a terrain.

### v4

@@ 2 Related Work
Recent progress in legged-robot motor skills follows two complementary streams. Reward-based methods use model-free RL in simulation, shaping behavior with handcrafted objectives that mix task terms (e.g., velocity tracking) and motion-naturalness regularizers; thanks to massive parallel physics engines [11, 12], this paradigm has produced agile locomotion on quadrupeds and humanoids without motion data. However, each new behavior demands tuning of user-defined rewards and environment scripting [13, 14, 15, 4, 6, 9, 16, 17]. Data-driven methods instead imitate reference motion, originally MoCap clips or monocular video, training a simulated character to track them and porting the idea to robots [1, 3, 18, 19, 7, 8]. For example, recent work [20, 21] frames legged locomotion as a next- token prediction task and pre-trains a policy on human data in kinematic space, showing strong performance. While imitation bypasses reward engineering [5], existing works typically assume flat ground or manually designed setups, limiting context-aware whole-body control; even animation systems that model human-scene contact rely on instrumented MoCap stages and thus lack scalability [22, 23]. Our system conditions on visual observations, the local height‑maps, and learns environment‑aware skills such as stair‑climbing and chair‑sitting directly from monocular RGB videos. Joint 4D human–scene reconstruction provides physically consistent reference motions, which RL distills into policies that transfer to a real humanoid (Table 1).
@@ 3 Real-to-Sim Data Acquisition
Our real-to-sim pipeline proceeds as summarized in Figure 2. We extract per-frame human poses and a raw scene point cloud from the input video (Sec. 3.1); jointly optimize them to obtain metrically aligned human trajectories and scene geometry (Sec. 3.2); apply gravity alignment and convert the filtered point cloud into a lightweight mesh (Sec. 3.3); and retarget the refined trajectories to the humanoid under joint-limit, contact, and collision constraints (Sec. A.3). The resulting motion–mesh pairs are ready for policy learning in Sec. 4.
@@ 4 Policy Learning
Stage 1: MoCap Pre-Training. MoCap pre-training lets a policy learn challenging skills from noisy video reconstruction while keeping hand-crafted priors to a minimum and bridging the human-to-robot embodiment gap. Earlier work tackles this either by sampling start poses with multi-agent RL [3] or by having a privileged simulator imitate the motion [5]. Radosavovic et al. [21] and Singh et al. [56] instead employed a form of kinematic pre‑training on human data. We adopt a simpler yet effective strategy: first pre‑train the policy on MoCap trajectories, then fine‑tune it on our reconstructed video data—both stages use reinforcement learning in a physics simulator. Even the MoCap‑only policy can be deployed directly on the real robot as shown in Fig 9. We used LAFAN motion capture data [57] retargeted to Unitree G1. For the pretrained policies, the conditioning the policy receives is the target joint angles, target root roll/pitch, and desired root direction.
@@ 5.1 Reconstruction and Data
We compare our method against baselines following the standard evaluation protocol [60, 47]. As summarized in Table 2, our method consistently achieves the best performance, outperforming prior work in both human trajectory accuracy (WA/W-MPJPE) and scene geometry (Chamfer Distance). Refer to Sec. A.4 for more evaluation details.
@@ 5.2 Real-world Deployment
Setup. We deploy our controller on a 23-DoF Unitree G1 humanoid and run it onboard at 50 Hz. Following [61], we set relatively low joint gains, $K_{p}=75$, to avoid excessively fast or overly stiff behaviour—which helps to avoid excessively violent contact when the robot makes heavy contact with objects such as chairs or stairs. Height-maps are computed in real time using Fast-lio2 [62] and probabilistic terrain mapping [63, 17]. We feed joystick targets from a human operator. Including policy running, all operations are run onboard. We found two critical ingredients for successful motion deployment through iterative sim-to-real trials: (i) relaxing the episode-termination tolerances with respect to the reference motion, and (ii) injecting realistic physics perturbations during training. Full details are given in Sec. C.
@@ Acknowledgments
We thank Brent Yi for his guidance with the excellent 3D visualization tool we use, Viser. We are grateful to Ritvik Singh, Jason Liu, Ankur Handa, Ilija Radosavovic, Himanshu Gaurav-Singh, Haven Feng, and Kevin Zakka for helpful advice and discussions during the paper. We thank Lea Müller for helpful discussions at the start of the project. We thank Zhizheng Liu for helpful suggestions on evaluating human and scene reconstruction. We thank Moji Shi and Huayi Wang for advice on using LiDAR heightmap input. We thank Eric Xu, Matthew Liu, Hayeon Jeong, Hyunjoo Lee, Jihoon Choi, Tyler Bonnen, and Yao Tang for their help in capturing and featuring in the video clips used in this project. We acknowledge support from the BAIR Humanoid Intelligence Center. Chung Min Kim is supported by the NSF Research Fellowship Program, Grant DGE 2146752. Pieter Abbeel holds concurrent appointments as a Professor at UC Berkeley and as an Amazon Scholar. This paper describes work performed at UC Berkeley and is not associated with Amazon. This project was funded in part by NSF: CNS-2235013, IARPA DOI/IBC No. 140D0423C0035, ONR MURI awards N00014-21-1-2801 and W911NF-23-1-0277, and Bakar fellows.
@@ A.2 Objectives for Joint Human–Scene Reconstruction
As shown in Figure 7(a), the coarse initialization produces inaccurate human trajectories and an incorrect scene scale. Our optimization procedure refines these estimates by jointly optimizing the human’s global translations ($\gamma^{1:T}$), global orientations ($\phi^{1:T}$), local poses ($\theta^{1:T}$), and the scene point cloud scale ($\alpha$). The optimization aligns the human and world as shown in Figure 7(b), and aids in generating simulation-ready data.
@@ B.2 Observations
Table 3 details the actor and critic observations. We note that, unlike works in graphics [1, 3], the actor gets all unprivileged observations that are all available on the real robot (see Sec. 5). Furthermore, the target root reference during train time is fed into the global frame to ensure it is aligned with the terrain. Thus, unlike [5, 8], we are able to train policies using a global root reference, allowing our policies to correctly imitate references in the world frame over long horizons, which is critical for correctly imitating motions tied to a terrain.